OpenStudy (anonymous):

Which ratio is equivalent to 18/25?

OpenStudy (anonymous):

18/25=36/50

OpenStudy (anonymous):

A. 5/8 B. 2/8 C. 10/12 D. 3/4

OpenStudy (anonymous):

18/25=.72

OpenStudy (mathmate):

@Lindsay_552 are you sure you have the choices for the problem above?

OpenStudy (anonymous):

D id correct as

OpenStudy (anonymous):

18/25 = .72

OpenStudy (anonymous):

as it would be 18/25 25 is greater then 18 so add a zero after 8

OpenStudy (anonymous):

so 180/25

OpenStudy (anonymous):

okay thanks everyone :)

OpenStudy (anonymous):

non of the above is an equivalent but the 3/4 closest